For 270.0 mL of a buffer solution that is 0.245 M in HCHO2 and 0.305 M in KCHO2, calculate the initial pH and the final pH after adding 0.005 mol of NaOH.
Answer #1

Concentration of HCHO2 = 0.245M

Concentration of KCHO2 = 0.305M

Volume of the buffer solution = 270.0 mL = 0.270L

number of moles = Molarity x volume of the solution in L

number of moles of HCHO2 = 0.245M x 0.270L = 0.06615 moles

number of moles of KCHO2 = 0.305M x 0.270L = 0.08235 moles

Ka of HCHO2 = 1.8 x10^-4

Ka = 1.8 x10^-4

-log(Ka) = -log( 1.8 x10^-4)

PKa = 3.74

At initial

PH = PKa + log[salt]/[acid]

PH = Pka + log[KCHO2]/[HCHO2]

PH = 3.74 + log( 0.08235/0.06615)

PH = 3.83

after addition of NaOH

number of moles of NaOH = 0.005 moles

after addition of NaOH

number of moles of HCHO2 = 0.06615 - 0.005 = 0.06115 moles

number of moles of KCHO2 = 0.08235 + 0.005 = 0.08735 moles

PH = Pka + log[salt]/[acid]

PH = 3.74 + log( 0.08735/0.06115)

PH = 3.89